VEGAS Team Meeting

Attendees: Amanda, Anish, Paul, Ray, Joe, Marty, Richard, Mark, Randy

Agenda:

  1. DONE Is OGP calibration implemented? Is ADC calibration in general complete?
  2. DONE Are stand-along ADC calibration routines ready to send to DIBAS. What about documentation to accompany them?
  3. DONE Mode 10-19 (ghost frequencies issue). Is Randy making any progress?
  4. DONE Blanking of “last switched integration” – is it implemented?
  5. DONE Observer documentation
  6. DONE Use of test time (including overnights)
  7. DONE Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues

Meeting notes:

  1. Is OGP calibration implemented? [No] Is ADC calibration in general complete? [No]
    • After MMC changes were promoted, continuum data taken in mode #1 looks stable - map looked good.
    • some warnings need to be investigated
    • Monitor the MMC part
    • OGP work remains
      • after the work is completed, possibly move rest frequency back to center of band?
      • Manager will look at stored values and reload for new observations
      • warning if memory vs. stored exceed a preset valus
  2. Are stand-along ADC calibration routines ready to send to DIBAS. What about documentation to accompany them?
    • Developed Python code will be ok for both platforms, but not ready yet
    • Documentation is up to date with progress.
  3. Mode 10-19 (ghost frequencies issue). Is Randy making any progress?
    • version 142 - still has issue
    • looking at two 'halves' of firmware and running simulations trying to reproduce problem - no success
    • Joe investigating ability to reproduce problem in just CPU with artificial packets - to be picked up after LO-1 work
    • Continue investigation
  4. Blanking of “last switched integration” – is it implemented?
    • Issue with LO-1, scan coordinator drives VEGAS to a value that exceeds LO-1 period.
    • Last integration that exceeds scan length is blanked is a solution that has been coded, but needs testing and deployment.
  5. Observer documentation
    • Upon investigation of existing documentation for a GBT Memo.
      • Proposer's guide data requires a lot of changes to make sure the baseline of information in proposals is sufficient to judge proposals
      • Amanda will take lead on edits
      • Observer's Guide will need similar updates as well
  6. Use of test time (including overnights)
    • Tuesday (inc overnight)
    • Wednesday (inc. overnights)
      • Unit tests as needed by Paul (coordinated w/GBT Operator)
    • Thursday
  7. Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues

Other

  1. Please make sure to properly code your timesheets for the current time period and going forward.
    • Software engineers should be using a GBT ops software CM&E account - Account 893252
    • Electronic engineers should be using a GBT ops electronics CM&E account - Account 432510
    • Commissioning activities should be getting charged to hardware support and testing - Account 432510
    • Scientists assisting with shared-risk commissioning should be charging an observer support account
      • 432540 for the most part. Accounts 142020 (GB-based staff) and 142010 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-06-10
Topic revision: r1 - 2014-06-10, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback